# RIs-Calib: An Open-Source Spatiotemporal Calibrator for Multiple 3D Radars and IMUs Based on Continuous-Time Estimation

### v1

@@ III-B1 Rotation B-spline Initialization
We first perform a rotation-only B-spline fitting based on the raw angular velocity measurements of IMUs to recover the rotation B-spline, where the extrinsic rotations of IMUs as by-products could be initialized simultaneously. Specifically, the least-squares problem could be described as:
@@ III-B2 Extrinsics and Gravity Initialization
After the rotation B-spline and extrinsic rotations of IMUs are initialized, we move on to initialize extrinsic translations of IMUs, extrinsics of radars, and gravity. Consider that the $j$-th radar takes one measurement of a static target at time $\tau$, based on (4) and subjected to the static constraint ${{}^{b^{c}_{0}}\dot{\boldsymbol{p}}_{t}}\equiv\boldsymbol{0}_{3\times 1}$, we have:
@@
$\mathcal{V}_{\tau_{k},\tau_{k+1}}^{i,\prime}$ and $\mathcal{V}_{\tau_{k},\tau_{k+1}}^{i,\prime\prime}$ are quantities that can be obtained by numerical integration based on the recovered rotation B-spline, initialized extrinsic rotations of IMUs, and linear acceleration measurements:
@@ III-B3 Velocity B-spline Initialization
With the initialized gravity vector, we recover the velocity B-spline and refine the quantities obtained in the previous step based on the raw measurements from radars and accelerometers. Specifically, the least-squares problem could be described as:
@@
where ${{}^{b^{c}_{0}}{\boldsymbol{v}}_{b^{c}}(\tau)}$ is the velocity of the central IMU at time $\tau$, which can be interpolated from the velocity B-spline. Note that we use the velocity-level pre-integration residual again when organizing this least-squares problem. The difference from the previous step is that the velocities of the central IMU, i.e., ${{}^{b^{c}_{0}}{\boldsymbol{v}}_{b^{c}}(\tau_{k})}$ and ${{}^{b^{c}_{0}}{\boldsymbol{v}}_{b^{c}}(\tau_{k+1})}$ in $\mathcal{V}_{\tau_{k},\tau_{k+1}}^{j}$, are obtained by directly interpolating the velocity B-spline, rather than by the radar velocities from a pre-solved linear least-squares problem. At this point, the initialization procedure is completed. It should be noted that we do not initialize the temporal parameters and intrinsics of IMUs, and set them as zeros or identities, which has little negative effect on the follow-up optimization.
@@ III-C3 Center Residual
Since we introduce a virtual central IMU and maintain its B-splines in the estimator, the center residuals are required to ensure the system has a unique least-squares solution. Specifically, we construct three types of center residuals, i.e., the rotational center residual, the translational center residual, and the temporal center residual. The rotational and translational center residuals have been defined in (9) and (13) as $\boldsymbol{r}_{ctr}(er)$ and $\boldsymbol{r}_{ctr}(et)$, respectively. In terms of the temporal center residual, we define it as:
@@ IV Simulation
Figure 4: The convergence performance of spatiotemporal parameters and IMU Biases in different steps: ($i$) INIT(1): Rotation B-spline Initialization (see Section III-B1); ($ii$) INIT(2): Extrinsics and Gravity Initialization (see Section III-B2); ($iii$) INIT(2): Velocity B-spline Initialization (see Section III-B3); $(iv)$ BO(j): the $j$-th batch optimization (see Section III-C). The white dashed line in each sub-figure indicates the zero line. All spatiotemporal parameters are with respect to IMU-1.
@@ V Real-world Experiment
TABLE I: The spatiotemporal calibration results of RIs-Calib in Monte-Carlo real-world experiments
Scenario Sensor Suite Extrinsic Temporal
Rotation Error Translation Error Time Offset
$\delta{\boldsymbol{\theta}(r)}$ $(deg)$ $\delta{\boldsymbol{\theta}(p)}$ $(deg)$ $\delta{\boldsymbol{\theta}(y)}$ $(deg)$ $\delta{\boldsymbol{p}(x)}$ $(cm)$ $\delta{\boldsymbol{p}(y)}$ $(cm)$ $\delta{\boldsymbol{p}(z)}$ $(cm)$ $\tau$ $(ms)$
Optimize Both Spatial and Temporal Parameters
Seq. 1 IMU-1 / IMU-2 0.01$\pm$0.01 0.02$\pm$0.01 0.02$\pm$0.02 0.03$\pm$0.02 0.01$\pm$0.01 0.05$\pm$0.03 -38.27$\pm$0.44
IMU-1 / RAD-1 0.74$\pm$0.86 0.61$\pm$0.95 0.51$\pm$0.46 0.46$\pm$0.31 0.38$\pm$0.16 0.29$\pm$0.17 -117.85$\pm$0.80
IMU-1 / RAD-2 0.44$\pm$0.70 0.69$\pm$0.97 0.41$\pm$0.18 0.31$\pm$0.20 0.46$\pm$0.15 0.21$\pm$0.19 -115.74$\pm$0.91
Seq. 2 IMU-1 / IMU-2 0.01$\pm$0.01 0.01$\pm$0.01 0.02$\pm$0.01 0.01$\pm$0.01 0.02$\pm$0.01 0.03$\pm$0.01 -37.19$\pm$0.16
IMU-1 / RAD-1 0.40$\pm$0.69 0.38$\pm$0.24 0.30$\pm$0.15 0.33$\pm$0.08 0.21$\pm$0.24 0.19$\pm$0.23 -116.64$\pm$0.98
IMU-1 / RAD-2 0.17$\pm$0.20 0.24$\pm$0.16 0.18$\pm$0.23 0.18$\pm$0.19 0.26$\pm$0.21 0.18$\pm$0.07 -115.51$\pm$0.50
Seq. 3 IMU-1 / IMU-2 0.02$\pm$0.01 0.01$\pm$0.01 0.01$\pm$0.02 0.01$\pm$0.01 0.03$\pm$0.02 0.02$\pm$0.01 -41.91$\pm$0.33
IMU-1 / RAD-1 0.76$\pm$0.86 0.36$\pm$0.34 0.42$\pm$0.37 0.29$\pm$0.10 0.17$\pm$0.12 0.24$\pm$0.24 -116.60$\pm$0.70
IMU-1 / RAD-2 0.37$\pm$0.39 0.76$\pm$0.77 0.21$\pm$0.05 0.36$\pm$0.11 0.34$\pm$0.06 0.28$\pm$0.29 -116.34$\pm$0.89
Optimize Only Spatial Parameters and Fix Temporal Ones as Identities
Seq. 1 IMU-1 / IMU-2 0.19$\pm$0.28 0.15$\pm$0.19 0.20$\pm$0.26 0.09$\pm$0.12 0.10$\pm$0.14 0.12$\pm$0.15 -
IMU-1 / RAD-1 1.14$\pm$2.69 1.47$\pm$2.40 1.23$\pm$2.98 1.11$\pm$1.27 1.29$\pm$1.64 0.42$\pm$0.51 -
IMU-1 / RAD-2 1.76$\pm$2.43 1.03$\pm$1.38 1.97$\pm$2.90 0.39$\pm$0.48 1.42$\pm$1.74 0.24$\pm$0.39 -
Seq. 2 IMU-1 / IMU-2 0.10$\pm$0.08 0.09$\pm$0.13 0.09$\pm$0.12 0.07$\pm$0.08 0.06$\pm$0.08 0.05$\pm$0.05 -
IMU-1 / RAD-1 1.20$\pm$1.57 0.79$\pm$0.98 0.93$\pm$1.17 1.25$\pm$0.87 1.01$\pm$0.79 0.80$\pm$0.84 -
IMU-1 / RAD-2 1.46$\pm$2.35 0.08$\pm$0.13 1.08$\pm$1.27 0.97$\pm$0.71 0.25$\pm$0.37 0.38$\pm$0.48 -
Seq. 3 IMU-1 / IMU-2 0.07$\pm$0.16 0.13$\pm$0.22 0.08$\pm$0.14 0.21$\pm$0.19 0.22$\pm$0.26 0.14$\pm$0.36 -
IMU-1 / RAD-1 1.12$\pm$1.28 1.26$\pm$1.40 0.98$\pm$1.17 0.82$\pm$0.98 1.17$\pm$1.25 0.82$\pm$1.13 -
IMU-1 / RAD-2 0.51$\pm$0.99 0.91$\pm$1.21 0.83$\pm$1.08 0.67$\pm$0.79 1.12$\pm$1.34 0.70$\pm$0.68 -
@@ V-A Accuracy, Repeatability, and Convergence Evaluation
We first evaluated our proposed RIs-Calib on calibration accuracy and repeatability. Altogether three data sequences collected in different scenarios were used, and each one of them was segmented into several data pieces lasting $50\;s$ for the Monte-Carlo tests. Fig. 7 depicts the calibration results of extrinsic translations, all of which are expressed with respect to IMU-1. The virtual central IMU is at the center of IMU-1 and IMU-2 due to the introduced center residuals. Intuitively, RIs-Calib possesses good repeatability, especially for the extrinsic calibration of two IMUs, whose range is within $1\;mm$. This is due to the high-frequency and high-precision inertial measurements IMUs provide. In terms of radars, their repeatability is lower than IMUs’ and the range is within $2\;cm$, which is reasonable since the accuracy of raw radar measurements is poor, especially the position observation of targets. Table I further summarizes the error distribution for calibration results of spatiotemporal parameters. The ground truth of extrinsics is obtained by computer-aided design (CAD). As for time offsets, considering the difficulty in obtaining the ground truth, calibration results are given by means and standard deviations (STDs) of estimates. As expected, the calibration results for extrinsic rotations and translations of IMUs are with higher accuracy and repeatability than those of radars. The extrinsic errors of IMUs are less than $0.02\;deg$ and $0.05\;cm$ with STDs within $0.02\;deg$ and $0.02\;cm$. As for radars, errors are $0.45\;deg$ and $0.30\;cm$ in average, and STDs are within $0.90\;deg$ and $0.50\;cm$. In terms of temporal parameters, since the synchronization of the platform is implemented on the software layer rather than the hardware layer, the time offsets could have subtle distinctions among separate startups, i.e., for three independently collected sequences. But in general, the STDs for the calibrated time offsets of both IMUs and radars are less than $1\;ms$. These results demonstrate that RIs-Calib could calibrate spatiotemporal parameters with high accuracy and excellent repeatability for multi-radar multi-IMU suites.
Most calibration works, especially the discrete-time-based ones, only focus on extrinsic determination, not temporal determination, limiting their applicability on most roughly synchronized low-cost devices. To further explore the impact of temporal calibration on extrinsic determination, additional comparative experiments were performed, and the results are summarized in Table I. In this experiment, temporal parameters are fixed as identities (zeros) and not involved in continuous-time batch optimization. As can be seen, for a sensor suite whose time synchronization is not conducted or is imprecise, ignoring temporal determination in calibration would introduce ineluctable errors. The extrinsic errors of IMUs increase to $0.2\;deg$ and $0.5\;cm$ with large STDs. As for two radars with larger time offsets than IMUs, extrinsic errors increase to $1.5\;deg$ and $1.5\;cm$, which is not unacceptable in accurate spatiotemporal calibration. These results illustrate the necessity of temporal determination in a calibration framework, especially orienting to roughly or not time-synchronized sensor suites.
@@ ACKNOWLEDGMENT
Shuolong Chen: Conceptualisation, Methodology, Software, Original Draft. Shengyu Li: Investigation, Formal analysis, Review and Editing. Xingxing Li: Resources, Supervision, Review and Editing, Funding acquisition. Yuxuan Zhou and Shiwen Wang: Data curation, Review and Editing.

### v2

@@ III-B1 Rotation B-spline Initialization
We first perform a rotation-only B-spline fitting based on the raw angular velocity measurements of IMUs to recover the rotation B-spline, where the extrinsic rotations of IMUs as by-products could be initialized simultaneously. This is achieved by solving the following least-squares problem:
@@ III-B2 Extrinsics and Gravity Initialization
After initializing extrinsic rotations of IMUs and the rotation B-spline of the reference IMU, we move on to initialize extrinsic translations of IMUs, extrinsics of radars, and gravity. Consider that the $j$-th radar takes one measurement of a static target at time $\tau$, based on (4) and subjected to the static constraint ${{}^{b^{c}_{0}}\dot{\boldsymbol{p}}_{t}}\equiv\boldsymbol{0}_{3\times 1}$, we have:
@@
$\mathcal{V}_{\tau_{k},\tau_{k+1}}^{i,\prime}$ and $\mathcal{V}_{\tau_{k},\tau_{k+1}}^{i,\prime\prime}$ are quantities obtained by numerical integration based on raw linear acceleration measurements and recovered states in least-squares problem (9):
@@ III-B3 Velocity B-spline Initialization
With the initialized gravity vector, we recover the velocity B-spline and refine the quantities obtained in the previous step based on the raw measurements from radars and accelerometers. The corresponding least-squares problem could be expressed as:
@@
where ${{}^{b^{c}_{0}}{\boldsymbol{v}}_{b^{c}}(\tau)}$ is the velocity of the central IMU at time $\tau$, which can be interpolated from the velocity B-spline. Note that we use the velocity-level pre-integration residual again when organizing this least-squares problem. The difference from the previous step is that the velocities of the central IMU, i.e., ${{}^{b^{c}_{0}}{\boldsymbol{v}}_{b^{c}}(\tau_{k})}$ and ${{}^{b^{c}_{0}}{\boldsymbol{v}}_{b^{c}}(\tau_{k+1})}$ in $\mathcal{V}_{\tau_{k},\tau_{k+1}}^{j}$, are obtained by directly interpolating the velocity B-spline, rather than by the radar velocities from a pre-solved linear least-squares problem.
For the left states in the estimators, such as intrinsics of IMUs, they are set as zeros or identities in initialization. At this point, the initialization procedure is completed.
@@ III-C3 Center Residual
Since we introduce a virtual central IMU and maintain its B-splines in the estimator, the center residuals are required to ensure the system has a unique least-squares solution. Adhering to our previous work [9], we construct three types of center residuals, i.e., the rotational center residual, the translational center residual, and the temporal center residual. The rotational and translational center residuals have been defined in (9) and (13) as $\boldsymbol{r}_{ctr}(er)$ and $\boldsymbol{r}_{ctr}(et)$, respectively. In terms of the temporal center residual, we define it as:
@@ IV Simulation
Figure 4: The system state and convergence performance in different stages: ($i$) INIT(1): Rotation B-spline Initialization (see Section III-B1); ($ii$) INIT(2): Extrinsics and Gravity Initialization (see Section III-B2); ($iii$) INIT(2): Velocity B-spline Initialization (see Section III-B3); $(iv)$ BO(j): the $j$-th batch optimization (see Section III-C). The white dashed line in each sub-figure indicates the zero line. For better readers’ understanding, all spatiotemporal parameters are with respect to IMU-1, rather than with respect to the virtual central IMU maintained in the estimator.
@@ V Real-world Experiment
TABLE I: The spatiotemporal calibration results of RIs-Calib and x-RIO in Monte-Carlo real-world experiments
Method Data Pair Extrinsic Temporal
Rotation Error $(deg)$ Translation Error $(cm)$ Time Offset
$\delta{\boldsymbol{\theta}(r)}$ $\delta{\boldsymbol{\theta}(p)}$ $\delta{\boldsymbol{\theta}(y)}$ $\delta{\boldsymbol{p}(x)}$ $\delta{\boldsymbol{p}(y)}$ $\delta{\boldsymbol{p}(z)}$ $\tau$ $(ms)$
RIs-Calib (Ours) Seq. 1 $b^{1}\leftrightarrow b^{2}$ 0.01 $\pm$0.01 0.02 $\pm$0.01 0.02 $\pm$0.02 0.03 $\pm$0.02 0.01 $\pm$0.01 0.05 $\pm$0.03 -38.27$\pm$0.44
$b^{1}\leftrightarrow r^{1}$ 0.74 $\pm$0.86 0.61 $\pm$0.95 0.51 $\pm$0.46 0.46 $\pm$0.31 0.38 $\pm$0.16 0.29 $\pm$0.17 -117.85$\pm$0.80
$b^{1}\leftrightarrow r^{2}$ 0.44 $\pm$0.70 0.69 $\pm$0.97 0.41 $\pm$0.18 0.31 $\pm$0.20 0.46 $\pm$0.15 0.21 $\pm$0.19 -115.74$\pm$0.91
Seq. 2 $b^{1}\leftrightarrow b^{2}$ 0.01 $\pm$0.01 0.01 $\pm$0.01 0.02 $\pm$0.01 0.01 $\pm$0.01 0.02 $\pm$0.01 0.03 $\pm$0.01 -37.19$\pm$0.16
$b^{1}\leftrightarrow r^{1}$ 0.40 $\pm$0.69 0.38 $\pm$0.24 0.30 $\pm$0.15 0.33 $\pm$0.08 0.21 $\pm$0.24 0.19 $\pm$0.23 -116.64$\pm$0.98
$b^{1}\leftrightarrow r^{2}$ 0.17 $\pm$0.20 0.24 $\pm$0.16 0.18 $\pm$0.23 0.18 $\pm$0.19 0.26 $\pm$0.21 0.18 $\pm$0.07 -115.51$\pm$0.50
Seq. 3 $b^{1}\leftrightarrow b^{2}$ 0.02 $\pm$0.01 0.01 $\pm$0.01 0.01 $\pm$0.02 0.01 $\pm$0.01 0.03 $\pm$0.02 0.02 $\pm$0.01 -41.91$\pm$0.33
$b^{1}\leftrightarrow r^{1}$ 0.76 $\pm$0.86 0.36 $\pm$0.34 0.42 $\pm$0.37 0.29 $\pm$0.10 0.17 $\pm$0.12 0.24 $\pm$0.24 -116.60$\pm$0.70
$b^{1}\leftrightarrow r^{2}$ 0.37 $\pm$0.39 0.76 $\pm$0.77 0.21 $\pm$0.05 0.36 $\pm$0.11 0.34 $\pm$0.06 0.28 $\pm$0.29 -116.34$\pm$0.89
x-RIO [31] Seq. 1 $b^{1}\leftrightarrow b^{2}$ 0.23 $\pm$0.31 0.35 $\pm$0.29 0.23 $\pm$0.18 0.91 $\pm$0.96 0.87 $\pm$0.82 0.73 $\pm$0.80
$b^{1}\leftrightarrow r^{1}$ 0.68 $\pm$0.94 0.67 $\pm$0.88 0.61 $\pm$0.79 1.23 $\pm$1.64 1.35 $\pm$0.72 1.40 $\pm$1.23
$b^{1}\leftrightarrow r^{2}$ 0.72 $\pm$0.74 0.61 $\pm$0.76 0.73 $\pm$0.86 0.98 $\pm$0.92 1.17 $\pm$1.39 1.01 $\pm$1.38
Seq. 2 $b^{1}\leftrightarrow b^{2}$ 0.30 $\pm$0.24 0.19 $\pm$0.18 0.27 $\pm$0.21 0.87 $\pm$0.72 1.08 $\pm$0.91 0.92 $\pm$0.98
$b^{1}\leftrightarrow r^{1}$ 0.75 $\pm$0.80 0.58 $\pm$0.92 0.63 $\pm$0.72 1.52 $\pm$1.47 1.20 $\pm$1.11 1.33 $\pm$1.07
$b^{1}\leftrightarrow r^{2}$ 0.62 $\pm$0.95 0.57 $\pm$0.72 0.66 $\pm$0.69 1.03 $\pm$0.94 1.12 $\pm$0.96 1.44 $\pm$1.26
Seq. 3 $b^{1}\leftrightarrow b^{2}$ 0.29 $\pm$0.27 0.32 $\pm$0.28 0.36 $\pm$0.30 0.71 $\pm$0.65 0.88 $\pm$0.83 0.90 $\pm$0.87
$b^{1}\leftrightarrow r^{1}$ 0.69 $\pm$0.54 0.56 $\pm$0.58 0.74 $\pm$0.75 1.34 $\pm$1.28 1.25 $\pm$1.09 1.14 $\pm$0.99
$b^{1}\leftrightarrow r^{2}$ 0.63 $\pm$0.70 0.78 $\pm$0.61 0.55 $\pm$0.89 1.21 $\pm$1.01 1.32 $\pm$1.17 1.25 $\pm$1.04
    Extrinsic translations in $(cm)$, extrinsic Euler angles in $(deg)$, and time offsets in $(ms)$.
    Cells with darker colors indicate larger errors for the corresponding parameters.
@@ V-A Accuracy, Repeatability, and Convergence Evaluation
We first evaluated our proposed RIs-Calib on calibration accuracy and repeatability. Altogether three data sequences collected in different scenarios were used, and each one of them was segmented into several data pieces lasting $50\;s$ for the Monte-Carlo tests. Fig. 7 depicts the calibration results of extrinsic translations of RIs-Calib, all of which are expressed with respect to IMU-1. The virtual central IMU is at the center of IMU-1 and IMU-2 due to the introduced center residuals. Intuitively, RIs-Calib possesses good repeatability, especially for the extrinsic calibration of two IMUs, whose range is within $1\;mm$. This is due to the high-frequency and high-precision inertial measurements IMUs provide. In terms of radars, their repeatability is lower than IMUs’ and the range is within $2\;cm$, which is reasonable since the accuracy of raw radar measurements is poor, especially the position observation of targets.
The open-source radar-inertial odometry x-RIO [31] is then selected as a benchmark for calibration comparison. Since x-RIO only supports online radar-IMU extrinsic calibration but not time offset calibration, we synchronized the radar data using the time offset estimated by RIs-Calib prior to conducting the x-RIO calibration to ensure a fair comparison. Additionally, we used the extrinsic calibration results from RIs-Calib as the initial values for x-RIO, as x-RIO, being an odometry system with online calibration capability, requires well-initialized extrinsics to achieve convergence. Table I summarizes the calibration error distribution of spatiotemporal parameters for both RIs-Calib and x-RIO. Cells with darker colors indicate larger errors for the corresponding parameters (a deeper red signifies greater extrinsic rotational error, while a deeper purple indicates larger extrinsic translational error). The ground truth of extrinsics is obtained by computer-aided design (CAD). As for time offsets, considering the difficulty in obtaining the ground truth, calibration results are given by means and standard deviations (STDs) of estimates. As expected, the calibration results for extrinsic rotations and translations of IMUs are with higher accuracy and repeatability than those of radars, which holds for both RIs-Calib and x-RIO. In comparison, RIs-Calib demonstrates overall superiority over x-RIO in calibration accuracy, particularly evident in the calibration of extrinsic translations. The extrinsic errors of IMUs from RIs-Calib are less than $0.02\;deg$ and $0.05\;cm$ with STDs within $0.02\;deg$ and $0.02\;cm$. As for radars, errors are $0.45\;deg$ and $0.30\;cm$ in average, and STDs are within $0.90\;deg$ and $0.50\;cm$. In terms of temporal parameters, since the synchronization of the platform is implemented on the software layer rather than the hardware layer, the time offsets could have subtle distinctions among separate startups, i.e., for three independently collected sequences. But in general, the STDs of time offsets calibrated by RIs-Calib are less than $1\;ms$. These results demonstrate that RIs-Calib could calibrate spatiotemporal parameters with high accuracy and excellent repeatability for multi-radar multi-IMU suites.
@@
    Extrinsic translations in $(mm)$, extrinsic Euler angles in $(deg)$, and time offsets in $(ms)$.
    Cells with darker colors indicate larger errors for the corresponding parameters.
@@ ACKNOWLEDGMENT
Shuolong Chen: Conceptualisation, Methodology, Software, Original Draft. Xingxing Li: Resources, Supervision, Review and Editing, Funding acquisition. Shengyu Li: Investigation, Formal analysis, Review and Editing. Yuxuan Zhou and Shiwen Wang: Data curation, Review and Editing.